OpenStudy (anonymous):

To find the height of a flagpole, a surveyor moves 90 feet away from the base of the flagpole and then, with a transit 3 feet tall, measures the angle of elevation to the top of the flagpole to be 60°. What is the height of the flagpole?

OpenStudy (anonymous):

Let x = length of flag pole from the transitthen tan(60)=x/90 90*tan(60)=x 155.88=x rounded 156 = x Now, add the height of the transit: 156+3=159 feet Answer-->159 feet